# Data-Driven Room Acoustic Modeling Via Differentiable Feedback Delay Networks With Learnable Delay Lines

### v2

@@ 4.5 Loss Function
where $\lambda\in\mathbb{R}_{\geq 0}$. Similarly to [19], the loss is evaluated in the time domain, and, at each iteration, requires a forward pass through the discrete-time model defined by the current parameter estimates. In the following sections, we will analyze each of the terms in (6).
@@ 4.5.1 Energy Decay Curve Loss
It is worth noting that, whereas the EDC is typically expressed in dB, (8) is evaluated on a linear scale. The idea is that a linear loss emphasizes errors in the early portion of $\varepsilon[n]$, i.e., where discrepancies are more perceptually relevant [37], compared to a logarithmic loss that would put more focus on the reverberation tail.
@@ 4.5.2 Differentiable Normalized Echo Density Profile
In [21], Abel and Huang introduced the so-called normalized echo density profile (EDP) as a means to quantify reverberation echo density by analyzing consecutive frames of the reverberation impulse response. The EDP indicates the proportion of impulse response taps that fall above the local standard deviation. The resulting profile is normalized to a scale ranging from nearly zero, indicating a minimal presence of echoes, to around one, denoting a fully dense reverberation with Gaussian statistics [38].
@@
First, we notice that the indicator function $\mathds{1}\!\left\{\lvert h[\tau]\rvert>\sigma\right\}$ can be equivalently expressed as a Heaviside step function $\mathcal{H}\left(\lvert h[\tau]\rvert-\sigma\right)$. Then, we let $g(x)$ denote the Sigmoid function. We define the scaled Sigmoid function $g_{\kappa}(x)=g(\kappa x)$, where $\kappa\in\mathbb{R}_{\geq 0}$. Since
@@
In practice, we would like to choose a large value of $\kappa$ but not larger than what is needed. Notably, the need for a large scaling factor is not constant throughout the temporal evolution of a RIR. Early taps are typically sparse, and $\left(\lvert h[\tau]\rvert-\sigma_{n}\right)$ tends to fall within the saturating region of $g_{\kappa}(\cdot)$, even for lower values of $\kappa$. Conversely, in later portions of the RIR, $\kappa$ must take on very large values to contrast the fact that the amplitude of $\left(\lvert h[\tau]\rvert-\sigma_{n}\right)$ progressively decreases. For this reason, we introduce a time-varying scaling parameter, $\kappa_{n}=\xi n+\varrho$, where $\xi\in\mathbb{R}_{>0}$ and $\varrho\in\mathbb{R}_{\geq 0}$ are hyperparameters. Progressively increasing the scaling coefficient has the benefit of enhancing the gradient flow for early decay, while improving the EDP approximation for late reverberation. In general, a more principled definition for $\kappa_{n}$ could be devised, e.g., by tying it to the local statistics of the target RIR or its energy decay. In this work, however, we favor a simple and reproducible approach as it proved to work well in practice.
@@ 4.5.3 Soft EDP Loss
Despite the existing trade-off between vanishing gradients and goodness of fit, every operation involved in the computation of (13) is differentiable. This allows us to use the following EDP loss term as a regularizer during the FDN training
@@
4.5.4 Model Summary
Figure 2 summarizes the proposed differentiable FDN model, listing all the unconstrained trainable parameters (top), their dimension, the corresponding reparameterization (middle), and illustrating once again how, through (1), the FDN processes time-domain signals, including unit impulses $\delta[n]$ (bottom).
Figure 2: Summary of the proposed FDN model.
@@ 5 Evaluation
We evaluate the proposed method using real-world measured RIR from the 2016 MIT Acoustical Reverberation Scene Statistics Survey [39]. The MIT corpus contains single-channel environmental IRs of both open and closed spaces. Of the 271 IRs, we select three according to their reverberation time, which, across the dataset, ranges from a minimum of $0.06$ s to a maximum of $1.99$ s. We select three indoor environments, namely, (i) a small room ($T_{60}\approx 0.2$ s), (ii) a medium room ($T_{60}\approx 0.6$ s), and (iii) a larger room ($T_{60}\approx 1.2$ s). For reproducibility, the ID of the chosen RIRs is reported: (i) h214_Pizzeria_1txts, (ii) h270_Hallway_House_1txts, and (iii) h052_Gym_WeightRoom_3txts. The full IR Survey dataset is available online.333[Online] IR Survey dataset: https://mcdermottlab.mit.edu/Reverb/IR_Survey.html For the evaluation, all IRs are resampled to 16 kHz and scaled to unit norm.
@@ 5.1 Baseline Methods
From the overview presented in Section 3, it appears that no method in the literature is directly comparable with ours. In fact, existing automatic tuning approaches either focus on off-the-shelf reverb plug-ins [26, 12], limit the set of target parameters to just a few [13, 14], or augment the FDN topology with auxiliary frequency-dependent components [15, 16]. To the best of our knowledge, there is no state-of-the-art method addressing the simultaneous estimation of every FDN parameters in a purely data-driven fashion.
@@
The first is based on a classic method for homogeneous reverberation time control (HRTC), and involves choosing all FDN parameters but the absorption coefficients heuristically. For simplicity, we refer to this method as “HRTC baseline.”
@@
The third and final baseline makes use of a genetic algorithm (GA) inspired by [16] to optimize every FDN parameter except for the feedback matrix. We call this method “GA baseline.”
In the following sections, we will detail the three baseline methods one at a time.
@@ 5.1.1 HRTC Baseline
Whereas the absorption coefficients are given by (16), every other parameter in the HRTC baseline are determined by means of heuristics. We parameterize the feedback matrix as in (2), where $\mathbf{U}$ is a random orthogonal matrix. To ensure that the HRTC baseline is most comparable with the proposed method (Section 4.3) and the Colorless baseline (Section 5.1.2), we obtain $\mathbf{U}$ through the exponential mapping in (3). The scalar gain $d$ is set equal to the amplitude of the target RIR at the time index associated with the direct path. We use unity input gains, i.e., $\mathbf{b}=\bm{1}_{N}$, where $\bm{1}_{N}$ is a vector of $N=6$ ones. The output gains are chosen so that $\mathbf{c}=\frac{1}{N}\bm{1}_{N}$. As such, the dot product $\mathbf{c}^{T}\mathbf{s}[n]$ in (1) is equivalent to the arithmetic average of the outputs of the $N$ delay lines at time $n$. Finally, the delays $\bm{m}=[997,1153,1327,1559,1801,2099]$ consist of the logarithmically distributed prime numbers from Delay Set #1 in [18].
@@ 5.1.2 Colorless Baseline
Colorless reverberation [41] is defined as the acoustic quality of an artificial reverberation algorithm whose frequency response is flat, i.e., constant at all frequencies.
To achieve this, $\mathbf{A}$, $\mathbf{b}$, and $\mathbf{c}$ are iteratively updated via backpropagation using Adam [42] to minimize a reference-free [^4] With reference-free, we emphasize that, unlike our method, the loss function in [18] is computed solely on the FDN response, and does not consider a reference RIR as the target of the optimization process. loss function comprising two terms [18]. The first term encourages the magnitude of the sampled transfer function to be as close to one as possible. The second term penalizes IR sparsity in the time domain and avoids trivial solutions.
@@ 5.1.3 GA Baseline
In [16], Ibnyahya and Reiss proposed a multi-stage automatic tuning approach that combines genetic optimization [43] and analytical filter design [44]. The prototype FDN considered in [16] is equipped with attenuation filters $H_{i}(z)$ controlling the decay rate and a tone-correction filter $T(z)$ [45]. We depict the model architecture with $N=3$ in Figure 3, featuring the output filter and graphic equalizers for each delay line (cf. Figure 1). Our implementation uses $N=6$.
As in [16], we aim to optimize $\bm{m}$, $\mathbf{b}$, $\mathbf{c}$, and $d$, whereas $\mathbf{A}$ is fixed throughout the procedure. The GA is run for $50$ generations (i.e., ten times more than in [16]), each with a population of $50$ FDNs. Each FDN is therefore an individual characterized by $13N+$ 1 mutable parameters, namely, $\bm{m}$, $\mathbf{b}$, $\mathbf{c}$, and $d$. During the optimization, scalar gains are constrained to take values in $[-1,1]$. Similarly, delays are constrained to take values in the range of 200 µs to 64 ms. The attenuation filters are analytically determined according to the individual’s delay values in $\bm{m}$ and the desired octave-band reverberation times [44]. In turn, the output graphic EQ filter [46] is found based on the initial level of the desired octave-band EDCs. All individuals implement the same random orthogonal feedback matrix [47], which is not affected by genetic optimization. The fitness of each individual at every generation is assessed through the mean absolute error between the MFCCs of the target RIR and those of the FDN output [16].
@@ 5.2 Evaluation Metrics
As evaluation metrics, we select the $T_{20}$, $T_{30}$, and $T_{60}$, i.e., the reverberation time extrapolated considering the normalized IR energy decaying from $-5$ dB to $-25$ dB, $-35$ dB, and $-65$ dB, respectively. Ideally, these three metrics are the same if the energy decay exhibits a perfectly linear slope. In practice, this is often not the case, as it can be seen, e.g., in Figure 11. Hence, we believe that it is more informative to report all three of them, as together they provide a richer insight into the global behavior of the EDC as it approaches the $-60$ dB threshold. It is also worth pointing out that it is unclear whether the $T_{60}$ is entirely reliable in measuring the reverberation time of real-world RIRs due to the often non-negligible noise floor.
Table 1: Reverberation time, loss values, iteration indices, composite loss function weights, and average time per training step. Iterations denoted with $(0)$ indicate pre-training random initialization.
$T_{60}$ iter $\mathcal{L}$ $\mathcal{L}_{\text{EDC}}$ $\mathcal{L}_{\text{EDP}}$ $\lambda$ s/iter
Gym $1.225$ $(0)$ $1.2070$ $0.9876$ $0.2194$ $1$ $23.73$
$279$ $0.0201$ $0.0091$ $0.011$
Hallway $0.607$ $(0)$ $1.0205$ $0.9904$ $0.3013$ $0.1$ $11.68$
$340$ $0.0057$ $0.0052$ $0.0046$
Pizzeria $0.206$ $(0)$ $0.9589$ $0.9345$ $0.2443$ $0.1$ $4.36$
$583$ $0.0781$ $0.0766$ $0.0154$
Furthermore, we report the following ISO 3382 measures: Clarity ($C_{80}$), expressed in dB, Definition ($D_{50}$), expressed as a percentage, and Center time ($t_{s}$), expressed in ms. Defined $L_{\tau}:=\left\lceil\tau\cdot f_{\mathrm{s}}\cdot 10^{-3}\right\rceil$, these metrics are given by [48]
@@
For all metrics considered, we report the error with respect to the target values. Absolute deviations are denoted by $\Delta$ in Tables 4, 4, and 4.
Finally, it is worth pointing out that the EDPs shown in the following sections are obtained using the (non-differentiable) formulation given in (9), unless explicitly stated otherwise.
@@ 5.3 Parameter Initialization
We initialize the differentiable FDN parameters as follows. We let $\mathbf{\tilde{b}}^{(0)}\sim\mathcal{N}(\bm{0},\frac{1}{N}\mathbf{I}_{N})$, where $\mathbf{I}_{N}$ is the $N\times N$ identity matrix. We let $\mathbf{\tilde{c}}^{(0)}=\frac{1}{N}\bm{1}_{N}$, where $\bm{1}_{N}$ is a vector of $N$ ones. We set $\tilde{d}^{(0)}=1$. We initialize $\mathbf{\tilde{W}}^{(0)}$ and $\bm{\tilde{\Gamma}}^{(0)}$ so that $\mathbf{\tilde{W}}^{(0)}_{ij}\sim\mathcal{N}(0,\frac{1}{N})$ and $\tilde{\gamma}^{(0)}_{i}\sim\mathcal{N}(0,\frac{1}{N})$. We initialize $\tilde{\bm{m}}^{(0)}$ so that $\tilde{m}^{(0)}_{i}=\psi\tilde{m}_{i}^{\star}$ with $\tilde{m}_{i}^{\star}\sim\text{Beta}(\alpha,\beta)$, for $i=1,...,N$, where $\alpha\geq 1$ and $\beta>\alpha$. We empirically set $\psi=1024$, $\alpha=1.1$, and $\beta=6$ to ensure a maximum possible delay of $64$ ms (the same as in the GA baseline) and a mean value of about $10$ ms. We let the Sigmoid scaling term $\kappa_{n}$ increase linearly from $10^{2}$ to $10^{5}$ as $n=0,...,L_{T_{60}}-1$. Finally, we choose the hyperparameter $\lambda$ for each test case (see Table 1) in order to balance the two terms in (6).

### v3

@@ 4.5 Loss Function
where $\lambda\in\mathbb{R}_{\geq 0}$. Similarly to [19], the loss is evaluated in the time domain, and, at each iteration, requires a forward pass through the discrete-time model defined by the current parameter estimates. In the following sections, we analyze each of the terms in (8).
@@ 4.5.1 Energy Decay Curve Loss
It is worth noting that, whereas the EDC is typically expressed in dB, (10) is evaluated on a linear scale. The idea here is that a linear loss emphasizes errors in the early portion of $\varepsilon[n]$, i.e., where discrepancies are perceptually more relevant [38], compared to a logarithmic loss that would put more focus on the reverberation tail.
@@ 4.5.2 Differentiable Normalized Echo Density Profile
In [21], Abel and Huang introduced the so-called normalized Echo Density Profile (EDP) as a means to quantify reverberation echo density by analyzing consecutive frames of the reverberation impulse response. The EDP indicates the proportion of IR taps that fall above the local standard deviation. The resulting profile is normalized to a scale ranging from nearly zero, indicating a minimal presence of echoes, to around one, denoting a fully dense reverberation with Gaussian statistics [39].
@@
First, we notice that the indicator function $\mathds{1}\!\left\{\lvert h[\tau]\rvert>\sigma\right\}$ can be equivalently expressed as a Heaviside step function $\mathcal{H}\left(\lvert h[\tau]\rvert-\sigma\right)$. Then, we let $g(x)$ denote the Sigmoid function. We define the scaled Sigmoid function $g_{\kappa}(x)=g(\kappa x)$, where $\kappa\in\mathbb{R}_{>0}$. Since
@@
In practice, we would like to choose a large value of $\kappa$ but not larger than what is needed. Notably, the need for a large scaling factor is not constant throughout the temporal evolution of a RIR. Early taps are typically sparse, and $\left(\lvert h[\tau]\rvert-\sigma_{n}\right)$ tends to fall within the saturating region of $g_{\kappa}(\cdot)$, even for lower values of $\kappa$. Conversely, in later portions of the RIR, $\kappa$ must take on very large values to contrast the fact that the amplitude of $\left(\lvert h[\tau]\rvert-\sigma_{n}\right)$ progressively decreases. For this reason, we introduce a time-varying scaling parameter, $\kappa_{n}=\xi n+\varrho$, where $\xi\in\mathbb{R}_{>0}$ and $\varrho\in\mathbb{R}_{\geq 0}$ are hyperparameters. Progressively increasing the scaling coefficient has the benefit of enhancing the gradient flow for the early reflections, while improving the EDP approximation for late reverberation. In general, a more principled definition for $\kappa_{n}$ could be devised, e.g., by tying it to the local statistics of the target RIR or its energy decay. In this work, however, we favor a simple and reproducible approach as it proved to work well in practice.
@@ 4.5.3 Soft EDP Loss
Despite the trade-off between vanishing gradients and goodness of fit discussed in the previous section, every operation involved in the computation of (15) is differentiable almost everywhere. This allows us to use the following EDP loss term as a regularizer during the FDN training
@@ 5 Evaluation
We evaluate the proposed method using real-world measured RIRs from the 2016 MIT Acoustical Reverberation Scene Statistics Survey [40]. The MIT corpus contains single-channel environmental IRs of both open and closed spaces. Of the $271$ IRs, we select three according to their reverberation time, which, across the dataset, ranges from a minimum of $0.06$ s to a maximum of $1.99$ s. We select three indoor environments:444 Although the proposed parameter tuning method shares some similarities with neural network training, particularly in their use of backpropagation, differentiable FDNs require a dedicated optimization routine for each target RIR. When it comes to evaluation, this study thus focuses on a limited number of illustrative examples; this approach is consistent with white-box system identification literature while contrasting with the way deep learning models are typically evaluated, which, instead, involves large-scale training and test sets. (i) a small room ($T_{60}\approx 0.2$ s), (ii) a medium room ($T_{60}\approx 0.6$ s), and (iii) a larger room ($T_{60}\approx 1.2$ s). For reproducibility, the ID of the chosen RIRs is reported: (i) h214_Pizzeria_1txts, (ii) h270_Hallway_House_1txts, and (iii) h052_Gym_WeightRoom_3txts. The full IR Survey dataset is available online.555[Online] IR Survey dataset: https://mcdermottlab.mit.edu/Reverb/IR_Survey.html For the evaluation, all IRs are resampled to 16 kHz and scaled to unit norm.
@@ 5.1 Baseline Methods
From the overview presented in Section 3, it appears that no method in the literature is directly comparable with ours. In fact, existing automatic tuning approaches either focus on off-the-shelf reverb plug-ins [26, 12], limit the set of target parameters to just a few [13, 14], or augment the FDN topology with auxiliary frequency-dependent components [15, 16]. To the best of our knowledge, there is no state-of-the-art method addressing the simultaneous estimation of every parameter of a time-invariant frequency-independent FDN in a purely data-driven fashion.
@@
The first is based on a classic method for homogeneous reverberation time control (HRTC) and involves choosing all FDN parameters but the absorption coefficients heuristically. For simplicity, we refer to this method as “HRTC baseline.”
@@
The third and final baseline, inspired by [16], makes use of a genetic algorithm (GA) to optimize every FDN parameter except for the feedback matrix. We call this method “GA baseline.”
In the following sections, we detail the three baseline methods one at a time.
@@ 5.1.1 HRTC Baseline
Whereas the absorption coefficients are given by (18), every other parameter in the HRTC baseline are determined by means of heuristics. We parameterize the feedback matrix as in (2), where $\mathbf{U}$ is a random orthogonal matrix. To ensure that the HRTC baseline is most comparable with the proposed method (Section 4.3) and the Colorless baseline (Section 5.1.2), we obtain $\mathbf{U}$ through the exponential mapping in (3). The scalar gain $d$ is set equal to the amplitude of the target RIR at the time index associated with the direct path. We use unity input gains, i.e., $\mathbf{b}=\bm{1}_{N}$, where $\bm{1}_{N}$ is a vector of $N=6$ ones. The output gains are chosen so that $\mathbf{c}=\frac{1}{N}\bm{1}_{N}$. As such, the dot product $\mathbf{c}^{T}\mathbf{s}[n]$ in (1) is equivalent to the arithmetic average of the outputs of the $N$ delay lines at time $n$. Finally, the delays $\bm{m}=[997,1153,1327,1559,1801,2099]$ consist of logarithmically distributed prime numbers from Delay Set #1 in [18], and the corresponding non-differentiable integer delay lines are implemented via buffer readout.
@@ 5.1.2 Colorless Baseline
Colorless reverberation [42] is here defined as the acoustic quality of an artificial reverberation algorithm whose frequency response is flat, i.e., constant at all frequencies.
To achieve this, $\mathbf{A}$, $\mathbf{b}$, and $\mathbf{c}$ are iteratively updated via backpropagation using Adam [43] to minimize a reference-free [^6] With reference-free, we emphasize that, unlike our method, the loss function in [18] is computed solely on the FDN response, and does not consider a reference RIR as the target of the optimization process. loss function comprising two terms [18]. The first term encourages the magnitude of the sampled transfer function of each delay network channel to be as close to one as possible. The second term penalizes IR sparsity in the time domain and avoids trivial solutions.
@@ 5.1.3 GA Baseline
In [16], Ibnyahya and Reiss proposed a multi-stage automatic tuning approach that combines genetic optimization [44] and analytical filter design [45]. Adhering to well-established design principles [7], the prototype FDN considered in [16] is equipped with attenuation filters $H_{i}(z)$ that modify the frequency and total energy of the normal modes of the system’s response, and a tone-correction filter $T(z)$ [46] that modify the system’s power spectral density by imposing a desired magnitude frequency response. While we depict the model architecture with $N=3$ in Figure 3, our implementation uses $N=6$.
As in [16], we aim to optimize $\bm{m}$, $\mathbf{b}$, $\mathbf{c}$, and $d$, whereas $\mathbf{A}$ is fixed throughout the procedure. The GA is run for $50$ generations (i.e., ten times more than in [16]), each with a population of $50$ FDNs. Each FDN is therefore an individual characterized by $13N+$ 1 mutable parameters, namely, $\bm{m}$, $\mathbf{b}$, $\mathbf{c}$, and $d$. During the optimization, scalar gains are constrained to take values in $[-1,1]$. Similarly, delays are constrained to take values in the range of 200 µs to 64 ms.
The attenuation filters are analytically determined according to the individual’s delay values in $\bm{m}$ and the desired octave-band reverberation times [45]. In turn, the output graphic EQ filter [47] is found based on the initial level of the desired octave-band EDCs. All individuals implement the same random orthogonal feedback matrix [48], which is not affected by genetic optimization. The fitness of each individual at every generation is assessed through the mean absolute error between the MFCCs of the target RIR and those of the FDN output [16].
@@
Table 1: Reverberation time, iteration indices, loss values, and average time per training step in seconds (NVIDIA Tesla V100). Iterations denoted with $(0)$ indicate pre-training random initialization.
$T_{60}$ iter $\mathcal{L}$ $\mathcal{L}_{\text{EDC}}$ $\mathcal{L}_{\text{EDP}}$ time [s]
Gym $1.225$ $(0)$ $0.9959$ $0.9768$ $0.1909$ $41.86$
$935$ $0.0067$ $0.0058$ $0.0095$
Hallway $0.607$ $(0)$ $1.0068$ $0.9812$ $0.2560$ $20.76$
$796$ $0.0041$ $0.0034$ $0.0068$
Pizzeria $0.206$ $(0)$ $0.9254$ $0.9038$ $0.2161$ $5.71$
$992$ $0.0526$ $0.0501$ $0.0255$
@@ 5.2 Evaluation Metrics
As evaluation metrics, we select the $T_{20}$, $T_{30}$, and $T_{60}$, i.e., the reverberation time extrapolated considering the normalized IR energy decaying from $-5$ dB to $-25$ dB, $-35$ dB, and $-65$ dB, respectively. Ideally, these three metrics are the same if the EDC exhibits a perfectly linear slope. In practice, this is often not the case, as it can be seen, e.g., in Figure 4. Hence, we believe that it is more informative to report all three of them, as together they provide a richer insight into the global behavior of the EDC as it approaches the $-60$ dB threshold. It is also worth pointing out that it is unclear whether the $T_{60}$ is entirely reliable in measuring the reverberation time of real-world RIRs due to the often non-negligible noise floor.
Furthermore, we report the following ISO 3382 measures [49]: Clarity ($C_{80}$), expressed in dB, Definition ($D_{50}$), expressed as a percentage, and Center time ($t_{s}$), expressed in ms. Having defined $L_{\tau}:=\left\lceil\tau\cdot f_{\mathrm{s}}\cdot 10^{-3}\right\rceil$, these metrics are given by
@@
For each metric, we report the error with respect to the target values. Absolute deviations are denoted by $\Delta$ in Tables 2, 3, and 4.
Finally, it is worth pointing out that the EDPs shown in the following sections are obtained using the (non-differentiable) formulation given in (11) unless explicitly stated otherwise.
@@ 5.3 Parameter Initialization
We initialize the differentiable FDN parameters as follows. We let $\mathbf{\tilde{b}}^{(0)}\sim\mathcal{N}(\bm{0},\frac{1}{N}\mathbf{I}_{N})$, where $\mathbf{I}_{N}$ is the $N\times N$ identity matrix. We let $\mathbf{\tilde{c}}^{(0)}=\frac{1}{N}\bm{1}_{N}$, where $\bm{1}_{N}$ is a vector of $N$ ones. We set $\tilde{d}^{(0)}=1$. We initialize $\mathbf{\tilde{W}}^{(0)}$ and $\bm{\tilde{\Gamma}}^{(0)}$ so that $\mathbf{\tilde{W}}^{(0)}_{ij}\sim\mathcal{N}(0,\frac{1}{N})$ and $\tilde{\gamma}^{(0)}_{i}\sim\mathcal{N}(0,\frac{1}{N})$. We initialize $\tilde{\bm{m}}^{(0)}$ so that $\tilde{m}^{(0)}_{i}=\psi\tilde{m}_{i}^{\star}$ with $\tilde{m}_{i}^{\star}\sim\text{Beta}(\alpha,\beta)$, for $i=1,...,N$, where $\alpha\geq 1$ and $\beta>\alpha$. We empirically set $\psi=1024$, $\alpha=1.1$, and $\beta=6$ to ensure a maximum possible delay of $64$ ms (the same as in the GA baseline) and a mean value of about $10$ ms. We let the Sigmoid scaling term $\kappa_{n}$ increase linearly from $10^{2}$ to $10^{5}$ as $n=0,...,L_{T_{60}}-1$.
5.4 Implementation Details
We implement our differentiable model in Python using PyTorch. We define the FDN as a class inheriting from nn.Module. We thus define the unconstrained trainable parameters as instances of nn.Parameter. Our model operates at a sampling rate of 16 kHz. As a result, its memory footprint turns out to be contained, allowing us to train all FDNs considered in the present study on a single 16 GB NVIDIA Tesla V100 graphics card. [^7] Preliminary experiments carried out with increased computational resources indicate that results comparable with those reported in the present study can be obtained at a sampling rate of 48 kHz. We optimize the models for a maximum of $1000$ iterations using Adam [43] with a learning rate of $0.1$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and no weight decay. In all test cases, the EDP loss term is weighted by $\lambda=0.1$.
The average training time per iteration is reported in Table 1. At each step, just below 24% of the time is taken by the forward pass of the FDN, approximately 12% is spent computing the loss function, and just above 64% is spent backpropagating the gradients and updating the parameters. Notably, we observe that the computation time increases linearly with the estimated $T_{60}$. For each test case, we present the model with the lowest composite loss.
5.5 Test Case: Gym (h052)
We start by considering the Gym RIR (h052). As shown in Table 1, the best model is reached at iteration 935, after the loss has decreased by three orders of magnitude with respect to the initial value obtained with the random initialization described in Section 5.3. Here, we present a comparison between the target room acoustics (solid black line), the GA baseline [16] (dotted blue line), the HRTC baseline (dash-dotted green line), the Colorless baseline [18] (dash-dotted blue line), and, finally, the proposed differentiable FDN (dashed orange line).
Figure 4, 5, and 6 show that the proposed method is capable of closely matching the EDC, EDP, and envelope of the target RIR, respectively. Conversely, the baseline methods produce poorer results.
In Figure 4, we may notice that the EDC of the GA baseline deviates from that of the target RIR after just 100 ms and exhibits an overall steeper decay. The EDC of both the HRTC and Colorless baselines, instead, overshoot the target after the very first few ms before linearly decaying. It is interesting to notice, though, that HRTC approaches and matches the target EDC at around the estimated $T_{60}$, i.e., $1.225$ s, differently from the Colorless baseline, which had overshoot the target curve more.
In Figure 5, the EDP of the baseline methods indicate a scarce echo density in the first 250 ms, especially for the HRTC and Colorless methods. The output of the FDN obtained with GA, instead, becomes identically zero just after approximately 0.9 s. The ensuing EDP pathologies are confirmed by the IR depicted in Figure 6, where the baselines are shown to yield fewer, more prominent taps compared to the IR of the proposed FDN model.
Notably, Figure 6 also suggests that all methods estimate a larger $d$ than what would correctly render the direct sound. We attribute this phenomenon to an attempt at compensating the lack of a noise floor that, in real-life measurements, contributes to the total energy of the RIR. We argue that, in FDN models with tunable direct gain, offsetting this bias is naively achieved by increasing $d$.
Table 2 shows that the proposed method has an overall better performance in five out of the six reverberation metrics, with a $\Delta T_{20}$, $\Delta T_{30}$, $\Delta C_{80}$, $\Delta D_{50}$, and $\Delta t_{s}$ of 16.5 ms, 55.2 ms, 0.02 dB, 0.09%, and 180 µs, respectively. In particular, $T_{20}$, $T_{30}$, and $t_{s}$ are estimated with an error over one order of magnitude lower than those of the other methods. Likewise, the proposed FDN improves upon the baselines by two orders of magnitude as far as clarity $C_{80}$ and definition $D_{50}$ are concerned. On the contrary, the proposed approach yields and error of 90.2 ms when it comes to the $T_{60}$, and it is surpassed by HRTC and Colorless, whose $\Delta T_{60}$ is 6.3 ms and 35.7 ms, respectively. This, however, was largely expected given that, in both baseline methods, the parameter $\gamma$ is specifically designed to match the desired $T_{60}$ according to (18).